Jefferson Lab HALL A Logbook for November 02 - All EXCEPT Automatic Entries

  • Main INDEX

    Tuesday

  • 90260 11/19/02 06:00 Higinbotham Shift Summary
  • 90259 11/19/02 05:56 Monaghan SRC Run Summary
  • Monday

  • 90246 11/18/02 22:02 lkaufman Hall C crosstalk tests summary
  • 90244 11/18/02 16:17 R. Michaels Fac. Dev. to continue
  • 90238 11/18/02 11:42 romanip Moller Target broken wire is fixed
  • 90237 11/18/02 09:52 R. Michaels Fac. Dev. update
  • 90235 11/18/02 07:52 Reitz Owl Shift Summary
  • Sunday

  • 90234 11/17/02 18:17 Vahe update on status
  • 90233 11/17/02 16:03 Reitz plan for fac. dev.
  • 90232 11/17/02 15:15 Reitz update on status
  • 90231 11/17/02 12:46 R. Michaels status update
  • 90230 11/17/02 08:13 paschke/kauf Owl Shift Summary
  • Saturday

  • 90226 11/16/02 23:45 Moffit LH2 put to sleep
  • 90225 11/16/02 23:36 BM, RS, DA Final Harp scan
  • 90224 11/16/02 23:25 Moffit empty target position
  • 90223 11/16/02 23:23 BM, RS, DA Study spot size effect
  • 90222 11/16/02 23:02 Moffit Target to 4cm LH2
  • 90221 11/16/02 22:32 Moffit Target to 15cm LH2
  • 90220 11/16/02 22:23 BM,RS,DA Change beam spot size
  • 90219 11/16/02 21:44 suleiman Beam Current During 4 cm LH2
  • 90218 11/16/02 21:00 BM, RS, DA Harp Scan
  • 90217 11/16/02 20:48 Moffit empty target position
  • 90216 11/16/02 20:40 BM, RS, DA Parity Taerget Boiling-part4: 4 cm LH2
  • 90215 11/16/02 20:21 Moffit LH2 fan to 60Hz
  • 90214 11/16/02 20:20 Moffit LH2 Fan to 70Hz
  • 90213 11/16/02 20:01 suleiman Beam Current During LH2
  • 90212 11/16/02 19:35 Moffit LH2 Fan speed 60Hz
  • 90211 11/16/02 19:34 Moffit H2 Gas Panel Alarm. Low Pressure
  • 90210 11/16/02 19:27 Moffit LH2 Fan Speed to 29Hz
  • 90209 11/16/02 19:21 Moffit LH2 Fan Speed to 70Hz
  • 90208 11/16/02 19:14 Moffit LH2 Fan Speed 40Hz
  • 90207 11/16/02 18:48 Moffit Target in 4cm LH2
  • 90206 11/16/02 18:45 BM, RS, JH Parity Taerget Boiling-part3: 15 cm LH2
  • 90205 11/16/02 17:49 suleiman Beam Current During Thick Carbon
  • 90204 11/16/02 17:44 Moffit LH2 Fan Speed back to 60Hz
  • 90203 11/16/02 17:36 Moffit LH2 Fan to 70Hz
  • 90202 11/16/02 17:30 Moffit LH2 Target Fan to 40Hz
  • 90201 11/16/02 17:30 Moffit Target to LH2
  • 90200 11/16/02 16:24 BM, RS, JH Parity Taerget Boiling-part2: Thick C
  • 90199 11/16/02 15:38 Moffit LH2 Target Awakened
  • 90198 11/16/02 15:22 Moffit LD2 Target put to sleep
  • 90197 11/16/02 14:59 Moffit Target in Thick Carbon
  • 90196 11/16/02 14:52 BM, RS, JH Parity Taerget Boiling-part1: LD2
  • 90195 11/16/02 14:50 Moffit Target Fan to 60Hz
  • 90194 11/16/02 14:46 Moffit Target Fan at 70Hz
  • 90193 11/16/02 14:39 Moffit Target Fan to 60Hz
  • 90192 11/16/02 14:35 Moffit Target Fan Speed at 30Hz
  • 90191 11/16/02 14:29 Moffit Fan Speed to 40Hz
  • 90190 11/16/02 13:19 Suleiman Beam current during LD2 Boiling
  • 90189 11/16/02 12:36 Moffit bpm12 looks good
  • 90188 11/16/02 11:28 Moffit Target in 15cm LD2
  • 90187 11/16/02 11:27 BM,RS,JH Harp Scan
  • 90186 11/16/02 10:22 Adaq End of Run 3207
  • 90185 11/16/02 10:19 Moffit empty target position
  • 90184 11/16/02 10:03 Suleiman New Lumi on the Pivot
  • 90183 11/16/02 10:02 R. Michaels End experiment, Fac. Dev. Plan
  • 90182 11/16/02 10:01 R. Michaels L-arm VDC efficiency 85%
  • 90181 11/16/02 10:00 Moffit Unser Monitor
  • 90180 11/16/02 09:59 Moffit Unser Monitor
  • 90176 11/16/02 08:45 Moffit DACnoise Calibration
  • 90173 11/16/02 08:01 pawel Owl Shift Summary (Nov. 15 & 16)
  • 90168 11/16/02 05:41 M Jones rebooted target ioc
  • 90167 11/16/02 05:25 M Jones target weirdness
  • 90153 11/16/02 00:04 Wendy Swing Shift Summary
  • Friday

  • 90138 11/15/02 14:23 vasulk LD2 fan set to 60 Hz
  • 90136 11/15/02 11:55 vasulk LD2 Temperature stable
  • 90135 11/15/02 11:08 vasulk LD2 loop temp fluctuation
  • 90134 11/15/02 10:04 rom/moffit/k missing cable replaced, iocse9 (bpm12)
  • 90111 11/15/02 00:08 Ulmer Swing Shift Summary
  • Thursday

  • 90105 11/14/02 22:18 H. Ibrahim Target Problems
  • 90099 11/14/02 16:09 M Jones day shift summary
  • 90086 11/14/02 08:12 P. Ambrozewi Shift Summary
  • 90072 11/14/02 02:54 R. Michaels CODA config NOROC4
  • 90070 11/14/02 02:49 M Jones Fastbus crate for roc4 dead
  • 90068 11/14/02 01:57 M Jones harp scan
  • 90067 11/14/02 01:45 atrig Trigger-Download
  • 90062 11/14/02 00:05 H. Ibrahim Swing Shift Summary
  • Wednesday

  • 90058 11/13/02 23:11 H. Ibrahim Run Control Error Messages !
  • 90051 11/13/02 20:57 H. Ibrahim Injector Problems
  • 90049 11/13/02 18:22 H. Ibrahim Beam Problems
  • 90041 11/13/02 16:02 LeRose Shift Summary
  • 90037 11/13/02 13:49 LeRose BCM calibration
  • 90032 11/13/02 11:34 moffit Empty Target Position
  • 90031 11/13/02 10:27 LeRose Controlled Access
  • 90022 11/13/02 08:06 pawel Shift Summary
  • 90000 11/13/02 00:03 A. Deur Shift Summary
  • Tuesday

  • 89990 11/12/02 21:47 x. jiang left arm Lecroy screen
  • 89989 11/12/02 21:46 x. jiang left arm high-voltage
  • 89984 11/12/02 21:10 weinstei vdc ok
  • 89981 11/12/02 20:41 weinstei vdc
  • 89971 11/12/02 16:03 C. C. Chang Shift Summary
  • 89969 11/12/02 13:19 C. C. Chang Beam tuning
  • 89964 11/12/02 10:46 H. Ibrahim Target IOC rebooted
  • 89960 11/12/02 09:50 hibrahim Loop 2 JTV changed
  • 89956 11/12/02 08:04 atrig Trigger-Download
  • 89955 11/12/02 08:03 atrig Trigger-Download
  • 89952 11/12/02 08:00 R. Michaels owl shift summary
  • 89951 11/12/02 07:55 R. Michaels download trigger
  • 89947 11/12/02 05:56 R. Michaels cycle quads
  • 89941 11/12/02 04:42 R. Michaels current limited to 50 uA
  • 89938 11/12/02 04:22 R. Michaels beam arrives, did HARP scan
  • 89931 11/12/02 00:12 R. Michaels cosmics checkout ok
  • Monday

  • 89921 11/11/02 20:47 X. jiang beam permit
  • 89920 11/11/02 16:08 LeRose Shift Summary
  • 89918 11/11/02 10:36 Markowitz Target status
  • 89917 11/11/02 09:29 LeRose No Beam
  • 89914 11/11/02 00:44 A. Camsonne Target put to sleep
  • Sunday

  • 89907 11/10/02 12:22 R. Michaels new BTA
  • 89903 11/10/02 07:14 M Jones machine update: no beam until noon or end of shift
  • Saturday

  • 89901 11/09/02 20:56 punjabi beam status.
  • 89899 11/09/02 15:54 W. Boeglin Shift Summary
  • 89898 11/09/02 11:40 Adaq Vacuum Problem
  • 89893 11/09/02 07:59 kramer Shift Summary
  • 89877 11/09/02 00:04 A. Deur Shift Summary
  • Friday

  • 89860 11/08/02 16:08 W. Boeglin Shift Summary
  • 89847 11/08/02 08:04 Z.-E. Mezian Shift summary
  • 89845 11/08/02 07:45 M Jones rebooted target IOC
  • Thursday

  • 89828 11/07/02 23:36 A. Deur Cryotarget
  • 89826 11/07/02 23:15 Hyde-Wright cryotarget
  • 89820 11/07/02 20:53 R. Michaels T9 re-appears
  • 89811 11/07/02 17:06 Hyde-Wright DAQ deadtime
  • 89810 11/07/02 17:06 R. Michaels missing T9
  • 89809 11/07/02 16:50 Hyde-Wright DAQ
  • 89806 11/07/02 16:09 LeRose Shift Summary
  • 89801 11/07/02 15:57 LeRose Magnet Screen
  • 89800 11/07/02 15:55 atrig Trigger-Download
  • 89796 11/07/02 15:12 LeRose Magnet screen
  • 89790 11/07/02 13:32 suleiman Summary of June Lumi Studies
  • 89784 11/07/02 08:04 Bosted Shift Summary
  • 89775 11/07/02 05:03 atrig Trigger-Download
  • 89760 11/07/02 00:03 Reitz shift summary
  • Wednesday

  • 89740 11/06/02 16:01 W. Boeglin shift summary
  • 89730 11/06/02 08:40 Eric Beam Stability
  • 89729 11/06/02 08:37 Hyde-Wright gas
  • 89727 11/06/02 08:06 Bosted Shift Summary
  • 89709 11/06/02 00:11 Ulmer Swing Shift Summary
  • Tuesday

  • 89707 11/05/02 22:07 atrig Trigger-Download
  • 89701 11/05/02 20:33 Ulmer punch through
  • 89697 11/05/02 20:03 ulmer no end of run
  • 89695 11/05/02 19:05 Ulmer Collimator punch through
  • 89692 11/05/02 18:41 Ulmer Multiple end of runs.
  • 89686 11/05/02 16:12 W. Boeglin Shift Summary
  • 89684 11/05/02 16:00 solvignon End of run number error
  • 89671 11/05/02 08:00 R. Michaels owl shift summary
  • 89669 11/05/02 07:29 R. Michaels beam trip rate vs current
  • 89666 11/05/02 06:53 R. Michaels tools screen, flow rate alarm
  • 89658 11/05/02 04:08 last-shift last shift summary (late entry)
  • 89655 11/05/02 02:56 R. Michaels dhist / dataspy fixed
  • 89654 11/05/02 01:52 R. Michaels ET failed, restart CODA
  • 89649 11/05/02 00:49 adaq test
  • Monday

  • 89634 11/04/02 19:18 W. Boeglin beam spot size (not raster size)
  • 89633 11/04/02 19:13 Eric & Werne BPM Calibration
  • 89589 11/04/02 16:02 W. Boeglin Shift Summary
  • 89581 11/04/02 12:28 boeglin magnet screen for Q3_g40
  • 89579 11/04/02 11:04 atrig Trigger-Download
  • 89573 11/04/02 08:19 R. Michaels new BTA not ready for prime time
  • 89569 11/04/02 08:03 R. Michaels owl shift summary
  • 89568 11/04/02 07:54 A. Camsonne Target : Fluctuations of the target vacuum
  • 89567 11/04/02 07:36 R. Michaels detmap woes
  • 89559 11/04/02 03:58 A. Camsonne Target Moved to 15 cm LD2
  • 89556 11/04/02 03:32 atrig Trigger-Download
  • 89555 11/04/02 03:30 R. Michaels magnets Q3_g20, L-arm front camera problem
  • 89554 11/04/02 03:24 A. Camsonne Target moved to optics target
  • 89553 11/04/02 03:18 R. Michaels Kinematics Q3_g20
  • 89552 11/04/02 03:16 R. Michaels halog end-run DT fixed
  • 89548 11/04/02 02:01 R. Michaels network DAQ switch rebooting itself
  • 89547 11/04/02 01:54 R. Michaels SunOS halog fixed
  • 89546 11/04/02 01:44 Adaq test of SunOS halog
  • 89545 11/04/02 01:27 R. Michaels New BTA software
  • 89544 11/04/02 01:18 A. Camsonne Efficiency run(s)
  • 89537 11/04/02 00:39 R. Michaels second file for run 2914
  • Sunday

  • 89533 11/03/02 23:56 Bosted Shift Summary
  • 89524 11/03/02 19:38 Eric DAQ problems
  • 89520 11/03/02 19:19 atrig Trigger-Download
  • 89519 11/03/02 19:15 a-onl Trigger-Download
  • 89518 11/03/02 19:14 a-onl Trigger-Download
  • 89517 11/03/02 19:09 a-onl Trigger-Download
  • 89513 11/03/02 16:07 Ulmer Day Shift Summary
  • 89509 11/03/02 14:35 Ulmer Energy lock removed for test
  • 89500 11/03/02 09:18 ulmer Left collimator reading
  • 89497 11/03/02 08:06 Z.-E. Mezian Shift Summary
  • 89491 11/03/02 05:43 M Jones problem with harp
  • 89473 11/03/02 00:05 W. Boeglin Shift Summary
  • Saturday

  • 89464 11/02/02 16:04 Ulmer Day Shift Summary
  • 89461 11/02/02 10:31 ulmer tails in right arm ph_tg, th_tg
  • 89459 11/02/02 08:03 Z.-E. Mezian Shift Summary
  • 89458 11/02/02 07:02 M Jones changes to online espace
  • 89457 11/02/02 06:04 Adaq beam info in espace analysis
  • Friday

  • 89450 11/01/02 23:59 W. Boeglin Shift Summary
  • 89445 11/01/02 19:46 Boeglin no beam
  • 89440 11/01/02 16:02 Reitz Shift Summary
  • 89433 11/01/02 09:18 Higinbotham Injector Pockel Cell Change
  • 89427 11/01/02 08:04 Z.-E. Mezian shift summary
  • 89409 11/01/02 00:12 Boeglin Shift Summary